OpenStudy (anonymous):

Solve each inequality for questions 1–3. 1. m – 7 < 6 (1 point) m < –13 m < 1 m < –1 m < 13 2. n – 8 > 5 (1 point) n > 13 n > 3 n > –3 n > –13 3. p + 5 < 10 (1 point) p < –15 p < 5 p < 15 p < –5 4. The science class is taking a trip to the Science Center. There are 20 students and an unknown number of adult chaperones going. The bus holds at most 35 people. What is the greatest number of adults who can chaperone? (1 point) 5 10 15 75

OpenStudy (anonymous):

well for question 4, there are students and adults on a bus. so number of students + number of adults = number of people on the bus. right? so 20+ number of adults = 35 and now you just have to find the number of adults

OpenStudy (raden):

1). m-7 < 6 just added by 7 to both sides what u get

OpenStudy (anonymous):

Inequalities are solved like equations with a few rules. Some rules apply where you have to change the sign...< turns to > and vice versa. Like equations, you try to get the variable on one side and everything else on the other side. Rules : (1) adding or subtracting the same number on both sides does not change the sign. Even if both numbers are negative. (2) multiplying/ dividing a POSITIVE number on both sides does not change the sign. (3) multiplying / dividing by a NEGATIVE number on both sides CHANGES the inequality sign 1. m - 7 < 6 (add 7 to both sides of the < sign) m - 7 + 7 < 6 + 7 (combine like terms) m < 13 ===answer 2. n - 8 > 5 (add 5 to both sides to get n by itself) n - 8 + 8 > 5 + 8 n > 13 ===answer 3. p + 5 < 10 (subtract 5 from both sides to get p by itself) p + 5 - 5 < 10 - 5 p < 5 ===answer 4. s stands for students c stands for chaparones s + c < or equal to 35 s = 20 now substitute 20 in for s in the inequality... s + c < or = 35 20 + c < or = 35 (subtract 20 on both sides to get c by itself) 20 - 20 + c < or = 35 - 20 c < or = 35 - 20 c < or = 15 ===answer The reason I wrote this as an inequality is because it does not say there will be 35 people on the bus, it says " at most " 35 people. This means it is not a definite number. It means 35 or less people. This means it is = or <( less) then 35. I hope I could help you understand this better :)